Ticket Number: HEA-NOC/20080326-3 Ticket Status: UPDATE
Ticket Type: scheduled Resolver: HEA-NOC
Ticket Opened: 20080326 10:30 UTC Problem Start: 20080329 09:00 UTC
Ticket Update: 20080428 15:55 UTC+1 Problem End: 20080412 19:00 UTC+1
Site/line: Bluenet BH - KP, Bluenet BH - TCD, Bluenet BH-Citywest,
Bluenet DCU - KP, Bluenet TCD - KP, Bluenet TCD-DCU
In order to use DWDM equipment on the Dublin Core fibre ring, we need to
determine the fibre characteristics. This testing entails disconnecting
each section of the fibre ring and connecting testing equipment. Hence
resilience in the Dublin core ring will be lost while testing takes place.
Loss of resilience Saturday(12th of April) for DCU, TCD, Servecentric.
20080326 10:20 EK
20080331 14:09 EK Testing was completed from DCU-TCD and TCD-CWT on
Saturday the 29th of March. Sunday testing has been re-sheduled to
Saturday the 5th of April. There will be no resilience in the Dublin core
ring this Saturday the 5th of April.
20080404 14:16 EK Due to the contrator rescheduling, this work has been
moved from this Saturday to Saturday the 12th of April.
20080425 16:17 EK Work was completed on Saturday the 12th of April. This
uncovered some higher than expected losses on the fibre patches in
Servecentric. In order to identify the exact location of these losses
additional testing will be carried out this Monday the 28th of April.
Hence there will be intermittent loss of resilience in the Dublin core
ring on Monday from 9am to 5:30pm.
20080428 15:55 BB Testing of the DCU/Blanch stretch has commenced
This ticket can be monitored at http://www.hea.net/tickets/20080326-3
HEAnet Limited. Registered in Ireland, No. 275301.